# The super learner for time-to-event outcomes: A tutorial

### v1

@@ Abstract
Estimating risks or survival probabilities conditional on individual characteristics based on censored time-to-event data is a commonly faced task. This may be for the purpose of developing a prediction model or may be part of a wider estimation procedure, such as in causal inference. A challenge is that it is impossible to know at the outset which of a set of candidate models will provide the best predictions. The super learner is a powerful approach for finding the best model or combination of models (‘ensemble’) among a pre-specified set of candidate models or ‘learners’, which can include parametric and machine learning models. Super learners for time-to-event outcomes have been developed, but the literature is technical and a reader may find it challenging to gather together the full details of how these methods work and can be implemented. In this paper we provide a practical tutorial on super learner methods for time-to-event outcomes. An overview of the general steps involved in the super learner is given, followed by details of three specific implementations for time-to-event outcomes. We cover discrete-time and continuous-time versions of the super learner, as described by Polley and van der Laan (2011), Westling and others (2023) and Munch and Gerds (2024). We compare the properties of the methods and provide information on how they can be implemented in R. The methods are illustrated using an open access data set and R code is provided.
@@ 1 Introduction
A challenge in prediction modelling, and more generally in estimating any conditional risk, is that it is impossible to know at the outset which of a set of candidate models will provide the best fit, this typically being measured in terms of the out-of-sample predictive performance. Analysts are faced with the task of choosing from a suite of different time-to-event models. Time-to-event data are very often subject to right censoring of event times, and analysis methods need to accommodate this (Andersen and others, 2021). The semi-parametric Cox proportional hazards model is widely used to develop risk prediction models (Cox, 1972). Parametric models such as the Weibull model are also used. Extensions include using the Cox model in combination with the Lasso (Tibshirani, 1997), ridge and elastic net regression approaches and generalised additive models (GAMs) (Bender and others, 2018). Non-parametric and machine learning approaches for risk prediction that avoid parametric (or semi-parametric) modelling assumptions have been developed, for example random survival forests (Ishwaran and others, 2008). Different methods for developing a prediction model are often applied, with the resulting models then being compared in terms of their predictive performance, but this can be ad hoc. The super learner (SL) is a powerful approach for finding the best model or combination of models (‘ensemble’) among a pre-specified set of candidate models or learners, often referred to as the SL library (Polley and others, 2011). The non-ensemble SL selects the single best model and the ensemble (or standard) SL finds an optimally weighted combination of the set of candidate learners, which is guaranteed to perform as well as the best performing model included in the SL library. The non-ensemble SL is often referred to as the discrete SL, but we do not use this terminology in the current paper in order to avoid any confusion with a discrete-time implementation of the SL. The ensemble SL is a type of stacking algorithm, which have a longer history (Wolpert, 1992; Breiman, 1996). We focus on the SL because of its flexibility to include parametric, semi-parametric and non-parametric methods, including machine learning methods of different ‘architectures’, and because it has been shown to have the oracle property. The oracle property says that the SL will perform at least as well as the best performing individual model among the candidates, with this being true for both the ensemble and non-ensemble SLs. Unlike more ad hoc processes for identifying the best performing model from a set of candidate learners, which are often not pre-specified, the SL can be pre-specified as an approach to deriving a best-performing model. The derivation of a SL is based on minimizing a so-called loss function, which is a function describing the difference between estimates from the model and true (observed) values.
@@
The SL was originally proposed for continuous or binary outcomes, for which there are many models that could be included as candidate learners in the SL (van der Laan and others, 2007; Polley and others, 2011). The SL was extended for use with time-to-event outcomes by Polley and van der Laan (2011), who described a discrete-time approach. A discrete-time approach enables the use of candidate learners that are suitable for binary outcomes, but has the disadvantage that flexible estimation of the baseline hazard is challenging. More recently, two versions of the SL that accommodate continuous time have been proposed by Westling and others (2023) and Munch and Gerds (2024). By working in continuous time, these SLs can use candidate learners, including Cox regression and its extensions, parametric survival models, random survival forests, and others. The different versions of the SL also differ in terms of how they handle censoring and in terms of the loss function used. These SL methods for survival outcomes have been described in a technical way and some of them are reported within papers whose primary focus is on using the SL within a wider causal inference estimation procedure. An existing tutorial focused on the SL for continuous or binary outcomes (Naimi and Balzer, 2018) is available. The aim of this paper is to provide an accessible tutorial outlining how to use the SL for prediction of time-to-event outcomes.
@@
The paper is organised as follows. In Section 2 we define the set-up and notation and give an overview of the general steps of the SL. In Section 3 we outline the SL that uses a discretised time scale, as described by Polley and van der Laan (2011). Section 4 outlines two continuous-time versions of the SL, which do not require a discretisation of the time scale (Westling and others, 2023; Munch and Gerds, 2024). The practical implementation of the methods in R is discussed in Section 5, and in Section 6 we illustrate the use of the methods using the ‘Rotterdam’ data set, which is an openly accessible data set available in the survival package in R (Therneau and others, 2024). We conclude with a discussion in Section 7. To implement the methods we make use of R packages and Github repositories made available by the methods developers. We also provide simplified code that aims to provide a simpler illustration of the practical application of some of the methods. Code for reproducing the results in Section 5 is provided at https://github.com/ruthkeogh/superlearner_survival_tutorial.
@@ 2.2 Overview of super learner steps
Before outlining the details of specific implementations of the SL, we provide an overview of the general steps involved. We suppose that there are $p$ candidate learners for estimating the target estimand. We do not give details of different candidate models, but as noted earlier these could include a mixture of parametric, semi-parametric and machine learning methods. The set of candidate learners can include methods of the same type but with different hyper-parameters, for example random forests with different maximum tree depth. They can also include methods that may perform some variable selection, such as the Lasso.
The general steps involved in fitting a SL are as follows, with further details provided below:
@@
    Obtaining cross-validated estimates. Use model $j$ fitted using $D_{-k}$ to obtain predictions of the target estimand in validation set $D_{k}$, for $j=1,\ldots,p$ and $k=1,\ldots,K$. This gives cross-validated estimates of the survival probability $S(\tau|X=x)$ for each individual in $D$.
@@
    Estimating the mean expected loss. A loss function is a measure of the difference between predictions of the target estimand and the true value. The value of the loss function is calculated for each individual using each candidate learner $j=1,\ldots,p$ in each validation set $k=1,\ldots,K$. The expected loss in validation data set $D_{k}$ using candidate learner $j$ (based on model $j$ fitted using $D_{-k}$ in step 3) is denoted $L_{k,j}$. The mean expected loss across the $K$ validation data sets is calculated for each candidate learner, denoted $\bar{L}_{j}$, and this is sometimes referred to as the cross-validated risk for candidate learner $j$.
@@
    Deriving an ensemble. The non-ensemble SL stops following step 4 and selects the model with the smallest cross-validated risk. The ensemble SL uses further steps to derive an optimal combination or ensemble of the candidate learners. Broadly, this involves fitting a regression model with the observed outcome as the dependent variable and with the predictions from each model (obtained in step 3) as the explanatory variables, e.g. if there are $p=10$ candidate learners then there will be 10 explanatory variables.
@@
The cross-validation used in the SL is crucial to guarantee the oracle property, as well as to avoid over-optimism. In step 2, the continuous-time methods that we will consider (Westling and others, 2023; Munch and Gerds, 2024) also involve fitting models for the censoring distribution to each training set $D_{-k}$. For step 4, one example of a loss function is the squared difference between the observed outcome and estimated survival probability, weighted to account for censoring (Polley and van der Laan, 2011). Other loss functions will be discussed in the context of the different versions of the SL described below. See Sonabend and others (2024) for general discussion on loss functions for time-to-event outcomes. The cross-validated risk referred to in step 4 should not be confused with the risk of the event occurring, and we will use the alternative term mean expected loss to avoid confusion. In step 5 the regression model used to derive the ensemble should be done on a scale that is compatible with the loss function and the regression coefficients need to be constrained to be non-negative and to sum to 1. One example is to use non-negative least squares, resulting in non-negative coefficients that can be thought of as weights given to each candidate learner, with a weight of 0 meaning that the corresponding candidate learner does not contribute to the SL ensemble.
@@
For the ensemble SL the process results in the coefficients (or weights) for the candidate learners, whereas the non-ensemble SL results in the best single model. To then use the resulting SL to obtain an estimate of the target estimand $S(\tau|X=x)$ for a new individual we need to refit the necessary candidate learners on the full data $D$. For the non-ensemble SL we only need to refit the selected model, and for the ensemble SL we refit each of the candidate learners that is assigned a non-zero weight in step 5. The final ensemble SL combines these models fitted on the full data with the coefficients estimated in step 5.
@@
The non-ensemble SL is essentially a “standard” prediction model chosen via cross-validation. However, when following the SL algorithm, this prediction model is obtained through a principled and pre-specifiable process of choosing among many different candidate learners, which can be of different types and architectures, as well as including models of the same type but with different tuning parameter choices.
3 Discrete-time super learner: Polley and van der Laan (2011)
@@ 3.1 Discrete data set-up and notation
Polley and van der Laan (2011) described a SL for survival outcomes, which involves using a discretised time scale. This turns the problem of predicting a time-to-event outcome into that of predicting a binary outcome (the discrete-time event indicator) across a series of time points, using individuals still at risk at those time points, enabling use of any method suitable for binary outcomes in the set of candidate learners. The models are discrete-time hazard models. Time is discretised by dividing each individual’s follow-up into a series of discrete time periods up to and including the period in which they have the event or are censored. To simplify the notation of the methods described below we assume, without loss of generality, that follow-up time is discretised into periods of unit length, $[t,t+1)$ for $t=0,\ldots,\tau-1$. We let $\Delta(t)$ denote an individual’s status at the end of period $t$, with $\Delta(t)=0$ for periods in which the individual remains at risk at the end of the period (or if the period ends with censoring) and $\Delta(t)=1$ for the period in which the individual has the event. This results in a data set in counting process format with several rows per person (in other words a ‘long’ data set) - one row for each time period in which they are at risk. We denote the conditional discrete-time hazard at time $t$ by $Q(t|X)=\Pr(t\leq T<t+1|X,T\geq t)$. The conditional survival probability at time $\tau$ can be expressed as $S(\tau|X=x)=\prod_{t=0}^{\tau-1}1-Q(t|X=x)$. The conditional survival probability of interest can therefore be estimated based on estimates of the conditional discrete-time hazards. If $\tau$ is not one of the times in the discrete time grid (i.e. it is in between times on the grid) then $\tau$ is replaced by $\lfloor\tau\rfloor$ in the product.
@@ 3.2 SL steps 1-3: Cross-validation
For step 1 of the general SL algorithm outlined in Section 2.2 the data are divided into $K$ folds. This should be done on the basis of individuals rather than on the basis of rows in the long data set, so that all the rows of data for a given individual are in the same fold. For step 2 we need to fit each candidate learner for the conditional hazard $Q(t|X)$ as a function of $t$. The outcome in these models is $\Delta(t)$ (for $t=0,\ldots,\tau-1$) and the predictors are $t$ and $X$. The models are fitted on the long data set, i.e. pooled across times. candidate learners for $Q(t|X)$ could be any method for binary outcomes. In their illustrative application Polley and van der Laan (2011) included logistic regression models, random forests, generalised additive models (GAMs), polyclass, and a neural network. As noted earlier, the same method can be used with different hyper-parameters, and Polley and van der Laan (2011) used GAMs with different degrees of freedom and logistic regression models with and without interaction terms.
Fitting a model for $Q(t|X)$ requires consideration of how to incorporate time into the model, i.e. how to model the baseline discrete-time hazard as a function of time. Including separate indicators for each time period $t=0,\ldots,\tau-1$ could often be infeasible due to many time periods and/or too few events within some periods. Time could be included as a contionuous variable in the candidate learners, with different functions of time used as separate variables. As an alternative, Polley and van der Laan (2011) proposed a two-stage procedure which involves first fitting a logistic regression including covariates $t$ and $X$, and then using the fitted values from this as the offset term in a GAM including a flexible form for $t$. The degrees of freedom in the GAM used in the second stage is then a tuning parameter. The first stage model could alternatively be a random forest or a GAM, before using the resulting fitted values as the offset in a GAM including $t$.
After fitting candidate learner $j$ in each of the $K$ training sets $D_{-k}$ ($j=1,\ldots,p$; $k=1,\ldots,K$), the resulting learner is used to obtain an estimate of the conditional hazard $\hat{Q}_{j}(t|X=x_{i})$ for each individual in the validation set $D_{k}$ ($k=1,\ldots,K$). The cross-validated estimate of the survival probability based on candidate learner $j$ is obtained as $\hat{S}_{j}(\tau|X=x_{i})=\prod_{t=1}^{\tau}1-\hat{Q}_{j}(t|X=x_{i})$ for each person $i$ in $D$.

### v2

@@ Abstract
Estimating risks or survival probabilities conditional on individual characteristics based on censored time-to-event data is a commonly faced task. This may be for the purpose of developing a prediction model or may be part of a wider estimation procedure, such as in causal inference. A challenge is that it is impossible to know at the outset which of a set of candidate models will provide the best risk estimates. The super learner is a powerful approach for finding the best model or combination of models (‘ensemble’) among a pre-specified set of candidate models or ‘learners’, which can include both ‘statistical’ models (e.g. parametric, semi-parametric models) and ‘machine learning’ models. Super learners for time-to-event outcomes have been developed, but the literature is technical and the full details of how these methods work and can be implemented in practice have not previously been presented in an accessible format. In this paper we provide a practical tutorial on super learner methods for time-to-event outcomes. An overview of the general steps involved in the super learner is given, followed by details of three specific implementations for time-to-event outcomes. These include the originally proposed super learner, which involves using a discrete time scale, and two more recently proposed versions of the super learner for continuous-time data. We compare the properties of the methods and provide information on how they can be implemented in R. The methods are illustrated using an open access data set and R code is provided.
@@ 1 Introduction
A challenge in prediction modelling, and more generally in estimating any conditional risk, is that it is impossible to know at the outset which of a set of candidate models will provide the best fit, this typically being measured in terms of the out-of-sample predictive performance. Analysts are faced with the task of choosing from a suite of different time-to-event models. Time-to-event data are very often subject to right censoring of event times, and analysis methods need to accommodate this (Andersen and others, 2021). The semi-parametric Cox proportional hazards model is widely used to develop risk prediction models (Cox, 1972). Parametric models such as the Weibull model are also used. Extensions include using the Cox model in combination with the Lasso (Tibshirani, 1997), ridge and elastic net regression approaches and generalised additive models (GAMs) (Bender and others, 2018). Non-parametric and machine learning approaches for risk prediction that avoid parametric (or semi-parametric) modelling assumptions have been developed, for example random survival forests (Ishwaran and others, 2008). Different methods for developing a prediction model are often applied, with the resulting models then being compared in terms of their predictive performance, but this can be ad hoc. The super learner (SL) is a powerful approach for finding the best model or combination of models (‘ensemble’) among a pre-specified set of candidate models or learners, often referred to as the SL library (Polley and others, 2011). The non-ensemble SL selects the single best model and the ensemble (or standard) SL finds an optimally weighted combination of the set of candidate learners, which is guaranteed to perform as well as the best performing learner included in the SL library. The non-ensemble SL is often referred to as the discrete SL, but we do not use this terminology in the current paper in order to avoid any confusion with a discrete-time implementation of the SL. The ensemble SL is a type of stacking algorithm, which have a longer history (Wolpert, 1992; Breiman, 1996). We focus on the SL because of its flexibility to include parametric, semi-parametric and non-parametric methods, including machine learning methods of different ‘architectures’, and because it has been shown to have the oracle property. The oracle property says that the SL will perform at least as well as the best performing individual learner among the candidates, with this being true for both the ensemble and non-ensemble SLs. Unlike more ad hoc processes for identifying the best performing model from a set of candidates, which are often not pre-specified, the SL can be pre-specified as an approach to deriving a best-performing model. The derivation of a SL is based on minimizing a so-called loss function, which is a function describing the difference between estimates from the model and true (observed) values.
@@
The SL was originally proposed for continuous or binary outcomes, for which there are many models that could be included as candidate learners in the SL (van der Laan and others, 2007; Polley and others, 2011). The SL was extended for use with time-to-event outcomes by Polley and van der Laan (2011), who described a discrete-time approach. A discrete-time approach enables the use of candidate learners that are suitable for binary outcomes, but has the disadvantage that flexible estimation of the baseline hazard is challenging. More recently, two versions of the SL that accommodate continuous time have been proposed by Westling and others (2023) and Munch and Gerds (2024, 2025). By working in continuous time, these SLs can use candidate learners including Cox regression and its extensions, parametric survival models, random survival forests, and others. The different versions of the SL also differ in terms of how they handle censoring and in terms of the loss function used. SL methods for survival outcomes (Polley and van der Laan, 2011; Westling and others, 2023; Munch and Gerds, 2024, 2025) have been described in a technical way and some of them are reported within papers whose primary focus is on using the SL within a wider causal inference estimation procedure. An existing tutorial focused on the SL for continuous or binary outcomes is available (Naimi and Balzer, 2018), but does not extend to time-to-event outcomes. The aim of this paper is to provide an accessible tutorial outlining how to use the SL for prediction of time-to-event outcomes.
@@
The paper is organised as follows. In Section 2 we define the set-up and notation and give an overview of the general steps of the SL. In Section 3 we outline the SL that uses a discretised time scale, as described by Polley and van der Laan (2011). Section 4 outlines two continuous-time versions of the SL, which do not require a discretisation of the time scale (Westling and others, 2023; Munch and Gerds, 2025). The practical implementation of the methods in R is discussed in Section 5, and in Section 6 we illustrate the use of the methods using the ‘Rotterdam’ data set, which is an openly accessible data set available in the survival package in R (Therneau and others, 2024). We conclude with a discussion in Section 7. To implement the methods we make use of R packages and Github repositories made available by the methods developers. We also provide simplified code that aims to provide a simpler illustration of the practical application of some of the methods. Code for reproducing the results in Section 6 is provided at https://github.com/ruthkeogh/superlearner_survival_tutorial.
@@ 2.2 Overview of super learner steps
Before outlining the details of specific implementations of the SL, we provide an overview of the general steps involved. We suppose that there are $p$ candidate learners for estimating the target estimand. We do not give details of different candidate learners, but as noted earlier these could include a mixture of parametric, semi-parametric and machine learning methods. The set of candidate learners can include methods of the same type but with different hyper-parameters, for example random forests with different maximum tree depth. It can also include methods that may perform some variable selection, such as the Lasso.
The SL involves use of cross-validation to fit and assess the performance of different models, and in the ensemble version of the SL to derive the ensemble. The general steps involved in fitting a SL are as follows, with further details provided below:
@@
    Obtaining cross-validated estimates. Use candidate learner $j$ fitted using $D_{-k}$ to obtain predictions of the target estimand in validation set $D_{k}$, for $j=1,\ldots,p$ and $k=1,\ldots,K$. This gives cross-validated estimates of the survival probability $S(\tau|X=x)$ for each individual in $D$.
@@
    Estimating the mean expected loss. A loss function is a measure of the difference between predictions of the target estimand and the true value. The value of the loss function is calculated for each individual using each candidate learner $j=1,\ldots,p$ in each validation set $k=1,\ldots,K$. The expected loss in validation data set $D_{k}$ using candidate learner $j$ (based on candidate learner $j$ fitted using $D_{-k}$ in step 3) is denoted $L^{(j,k)}$. The mean expected loss across the $K$ validation data sets is calculated for each candidate learner, denoted $\bar{L}^{(j)}$, and this is sometimes referred to as the cross-validated risk for candidate learner $j$.
@@
    Deriving an ensemble. The non-ensemble SL stops following step 4 and selects the candidate learner with the smallest cross-validated risk. The ensemble SL uses further steps to derive an optimal combination or ensemble of the candidate learners. Broadly, this involves fitting a regression model with the observed outcome as the dependent variable and with the predictions from each candidate learner (obtained in step 3) as the explanatory variables, e.g. if there are $p=10$ candidate learners then there will be 10 explanatory variables.
@@
The cross-validation used in the SL is crucial to guarantee the oracle property, as well as to avoid over-optimism. In step 2, the continuous-time methods that we will consider (Westling and others, 2023; Munch and Gerds, 2025) also involve fitting models for the censoring distribution to each training set $D_{-k}$. For step 4, one example of a loss function is the squared difference between the observed outcome and estimated survival probability, weighted to account for censoring (Polley and van der Laan, 2011). Other loss functions will be discussed in the context of the different versions of the SL described below. See Sonabend and others (2024) for general discussion on loss functions for time-to-event outcomes. The cross-validated risk referred to in step 4 should not be confused with the risk of the event occurring, and we will use the alternative term mean expected loss to avoid confusion. In step 5 the regression model used to derive the ensemble should be done on a scale that is compatible with the loss function and the regression coefficients need to be constrained to be non-negative and to sum to 1. One example is to use non-negative least squares, resulting in non-negative coefficients that can be thought of as weights given to each candidate learner, with a weight of 0 meaning that the corresponding candidate learner does not contribute to the SL ensemble.
@@
For the ensemble SL the process results in the coefficients (or weights) for the candidate learners, whereas the non-ensemble SL results in the best single learner. To then use the resulting SL to obtain an estimate of the target estimand $S(\tau|X=x)$ for a new individual we need to refit the necessary candidate learners on the full data $D$. For the non-ensemble SL we only need to refit the selected learner, and for the ensemble SL we refit each of the candidate learners that is assigned a non-zero weight in step 5. The final ensemble SL combines estimates from these learners fitted on the full data with the coefficients estimated in step 5.
@@
The non-ensemble SL is essentially a “standard” prediction model chosen via cross-validation. However, when following the SL algorithm, this prediction model is obtained through a principled and pre-specifiable process of choosing among many different candidate learners, which can be of different types and architectures, as well as including learners of the same type but with different tuning parameter choices.
3 Discrete-time super learner
@@ 3.1 Discrete data set-up and notation
Polley and van der Laan (2011) described a SL for survival outcomes, which involves using a discretised time scale. Time is discretised by dividing each individual’s follow-up into a series of discrete time periods up to and including the period in which they have the event or are censored, with each period having an associated indicator of whether the individual had the event of interest in that period or remained event-free at the end of the period. This turns the problem of predicting a time-to-event outcome into that of predicting a binary outcome (the discrete-time event indicator) across a series of time points, using individuals still at risk at those time points, enabling use of any method suitable for binary outcomes in the set of candidate learners. The learners are then models for the discrete-time hazard. To simplify the notation of the methods described below we assume, without loss of generality, that follow-up time is discretised into periods of unit length, $[t,t+1)$ for $t=0,\ldots,\tau-1$. We let $\Delta(t)$ denote an individual’s status at the end of period $t$, with $\Delta(t)=0$ for periods in which the individual remains at risk at the end of the period (or if the period ends with censoring) and $\Delta(t)=1$ for the period in which the individual has the event. This results in a data set in counting process format with several rows per person (in other words a ‘long’ data set) - one row for each time period in which they are at risk. We denote the conditional discrete-time hazard at time $t$ by $Q(t|X)=\Pr(t\leq T<t+1|X,T\geq t)$. The conditional survival probability at time $\tau$ can be expressed as $S(\tau|X=x)=\prod_{t=0}^{\tau-1}1-Q(t|X=x)$. The conditional survival probability of interest can therefore be estimated based on estimates of the conditional discrete-time hazards. If $\tau$ is not one of the times in the discrete time grid (i.e. it is in between times on the grid) then $\tau$ is replaced by $\lfloor\tau\rfloor$ in the product.
@@ 3.2 SL steps 1-3: Cross-validation
For step 1 of the general SL algorithm outlined in Section 2.2 the data are divided into $K$ folds. This should be done on the basis of individuals rather than on the basis of rows in the long data set, so that all the rows of data for a given individual are in the same fold. For step 2 we need to fit each candidate learner for the conditional hazard $Q(t|X)$ as a function of $t$. The outcome in these learners is $\Delta(t)$ (for $t=0,\ldots,\tau-1$) and the predictors are $t$ and $X$. The learners are fitted on the long data set, i.e. pooled across times. Candidate learners for $Q(t|X)$ could be any method for binary outcomes. In their illustrative application Polley and van der Laan (2011) included logistic regression models, random forests, generalised additive models (GAMs), polyclass, and a neural network. As noted earlier, the same type of learner can be used with different hyper-parameters, and Polley and van der Laan (2011) used GAMs with different degrees of freedom and logistic regression models with and without interaction terms.
Fitting a model for $Q(t|X)$ requires consideration of how to incorporate time into the model, i.e. how to model the baseline discrete-time hazard as a function of time. Including separate indicators for each time period $t=0,\ldots,\tau-1$ could often be infeasible due to many time periods and/or too few events within some periods. Time could be included as a continuous variable in the candidate learners, with different functions of time used as separate variables. As an alternative, Polley and van der Laan (2011) proposed a two-stage procedure which involves first fitting a logistic regression including covariates $t$ (as a continuous variable) and $X$, and then using the fitted values from this as the offset term in a GAM including a flexible form for $t$. The degrees of freedom in the GAM used in the second stage is then a tuning parameter. The first stage model could alternatively be a random forest or a GAM, before using the resulting fitted values as the offset in a GAM including $t$ in the second stage.
After fitting candidate learner $j$ in each of the $K$ training sets $D_{-k}$ ($j=1,\ldots,p$; $k=1,\ldots,K$), the resulting learner is used to obtain an estimate of the conditional hazard $\hat{Q}_{j}(t|X=x_{i})$ for each individual $i$ in the validation set $D_{k}$ ($k=1,\ldots,K$). The cross-validated estimate of the survival probability based on candidate learner $j$ is obtained as $\hat{S}_{j}(\tau|X=x_{i})=\prod_{t=1}^{\tau}1-\hat{Q}_{j}(t|X=x_{i})$ for each person $i$ in $D$.